# PECAN: Personalizing Robot Behaviors through a Learned Canonical Space

### v1

@@ Abstract.
Robots should personalize how they perform tasks to match the needs of individual human users. Today’s robot achieve this personalization by asking for the human’s feedback in the task space. For example, an autonomous car might show the human two different ways to decelerate at stoplights, and ask the human which of these motions they prefer. This current approach to personalization is indirect: based on the behaviors the human selects (e.g., decelerating slowly), the robot tries to infer their underlying preference (e.g., defensive driving). By contrast, our paper develops a learning and interface-based approach that enables humans to directly indicate their desired style. We do this by learning an abstract, low-dimensional, and continuous canonical space from human demonstration data. Each point in the canonical space corresponds to a different style (e.g., defensive or aggressive driving), and users can directly personalize the robot’s behavior by simply clicking on a point. Given the human’s selection, the robot then decodes this canonical style across each task in the dataset — e.g., if the human selects a defensive style, the autonomous car personalizes its behavior to drive defensively when decelerating, passing other cars, or merging onto highways. We refer to our resulting approach as PECAN: Personalizing Robot Behaviors through a Learned Canonical Space. Our simulations and user studies suggest that humans prefer using PECAN to directly personalize robot behavior (particularly when those users become familiar with PECAN), and that users find the learned canonical space to be intuitive and consistent. See videos here: https://youtu.be/wRJpyr23PKI
@@
copyright: none
@@ 2. Related Work
Preference Learning. Humans can personalize a robot by providing feedback about its behaviors. This includes demonstrating the desired motion (Nikolaidis et al., 2015; Ravichandar et al., 2020), correcting the robot’s actions (Li et al., 2021; Spencer et al., 2022; Losey et al., 2022), selecting their preferred trajectory from options presented by the robot (Sadigh et al., 2017; Wilde et al., 2020; Zhan et al., 2021), or some combination of these feedback types (Jain et al., 2015; Munzer et al., 2017; Bıyık et al., 2022a; Mehta and Losey, 2023). These works parameterize the robot’s behavior with user-defined features, such as the speed of an autonomous car or its distance from other cars in a driving scenario. The parameters or weights for these features are then estimated based on the human’s feedback. Together, the weights and features determine the robot’s style (i.e., how it performs the given task) (Rosbach et al., 2019).
Specifying the correct set of features for modeling complex styles can be challenging (Bobu et al., 2024). To overcome this, previous research has also proposed deep reinforcement learning from human feedback (Christiano et al., 2017; Hejna III and Sadigh, 2023), which bypasses the need to specify such features. However, this approach is often time-consuming, as users may have to provide feedback hundreds or even thousands of times throughout the learning process.
@@
Learning a Canonical Style Space. Most relevant to our approach are methods like (Lynch et al., 2020; Osa and Ikemoto, 2020; Xihan et al., 2022; Śmieja et al., 2020) that learn a canonical representation of robot skills or styles over several tasks.
@@
In (Lynch et al., 2020), the robot embeds action sequences into a latent space of task-agnostic skills. The robot then executes these skills on a continuous range of tasks that are specified by their start and end states. When the set of tasks is finite, (Osa and Ikemoto, 2020) learns a latent space comprised of discrete and continuous portions. The discrete variables capture the tasks and the continuous variables capture the latent styles. Both these methods employ Variational Autoencoders (VAEs) to train the latent spaces in an unsupervised manner. However, (Lynch et al., 2020) assumes that the tasks are specified by the user, while (Osa and Ikemoto, 2020) does not guarantee that the respective latent spaces will exclusively capture the tasks and styles present in the data.
To address this, (Xihan et al., 2022) utilizes a small set of labels for trajectories that belong to the same task and for trajectories that correspond to the same skill. These labels are used to train Gated VAEs (Vowels et al., 2020) which encode the task and skill knowledge into separate latent spaces. In contrast, (Śmieja et al., 2020) encodes classes (i.e., tasks) and styles as a Gaussian mixture within a single continuous latent space. Each Gaussian represents a class and its variance captures the styles. As opposed to (Xihan et al., 2022), this method only requires the task labels for a subset of the inputs.
@@
A major limitation of all these methods is that the encoded styles are not consistent across tasks. The same latent value can produce different styles depending on the task — making the interface unintuitive for humans. Ideally, users should be able to select their preferred style with a single click and produce similar robot behaviors across all tasks.
@@
Similar to approaches like (Osa and Ikemoto, 2020), we utilize separate discrete and continuous latent spaces to encode the task and style information. The latent style space becomes our canonical space. To ensure that the canonical space is consistent across tasks, we use a small set of labels for trajectories having similar styles but in different tasks. Unlike (Xihan et al., 2022) and (Śmieja et al., 2020), our approach does not require any task labels. Instead, we capture the actual tasks and styles with their respective latent spaces by only using labels for trajectories with similar styles.
@@ 4.1. Separately Encoding Tasks and Styles
When this loss is minimized, it indirectly encourages the task and style encoders to capture sufficient representations of the trajectories in the dataset. However, this still does not guarantee that the representation captured in $Z{\tau}$ aligns with the actual tasks $\mathcal{T}$, or — along the same lines — that the representation in $Z_{\theta}$ aligns with the styles $\theta$.
To address this, we propose using a small set of labels $\mathcal{Y}$ for trajectories in the dataset. These labels identify trajectories from different tasks that share similar styles. In Section 4.3, we will introduce an additional loss function that leverages these labels to ensure that the representation captured in $Z_{\theta}$ aligns with the styles $\theta$. Further, we will show that by minimizing this loss together with $\mathcal{L}_{trajectory}$, we can also ensure that the latent tasks $z_{\tau}$ align with the actual tasks $\tau$.
@@ 4.2. Characteristics of a User-Friendly Canonical Space
    (2) $\theta(\phi(z_{\tau},z_{\theta}))\approx\theta(\phi(z_{\tau}^{\prime},z_{\theta}))\;\;\;\forall\;z_{\tau},z_{\tau}^{\prime}\in Z_{\tau}$
@@
  • Monotonicity. The styles should vary monotonically as we move from one point in the latent space to another. For example, decreasing the latent style value from $z_{\theta}=+1$ to $z_{\theta}=-1$ should gradually change the robot’s style from straight-line paths to increasingly exaggerated arm motions. Therefore, for all $z_{\theta},z_{\theta}^{\prime}\in Z_{\theta}$:
    (3) $(z_{\theta}-z_{\theta}^{\prime})(\theta(\phi(z_{\tau},z_{\theta}))-\theta(\phi(z_{\tau},z_{\theta}^{\prime})))\geq 0$
@@ 4.3. Semi-supervised Learning
Labeling Style Extremes. We first obtain a small set of labels $\mathcal{Y}$ for trajectories having similar styles but from different tasks. Specifically, we only obtain labels for trajectories that represent the extremes of the style range, e.g., the most or least exaggerated arm motions in Figure 2. We believe that labeling the extremes is easier than labeling intermediate styles. For instance, users find it difficult to distinguish between slightly different arm motions (Bıyık et al., 2019). Note that we do not ask users to specify the actual style of a trajectory. We only assume that trajectories with the same label have similar styles, and correspond to one of the extremes of the canonical space.
@@
We consider each $y_{i}\in\mathcal{Y}$ to be a separate class with one-hot labels $c(y_{i})$, where $\mathcal{Y}$ contains $m$ classes. At training time, we pass the labeled trajectories $\xi_{j}\in y_{i}$ through the style encoder $\psi_{\theta}$ to obtain their latent styles $z_{\theta,j}$. We then feed the latent styles into a classifier network $\Delta$ that maps each latent value to a $m$-dimensional vector of class probabilities $p_{j}=[p_{1},\ldots,p_{m}]$ such that $\sum_{k=1}^{m}p_{j,k}=1$ for any $p_{j}\in\mathcal{P}$.
@@
Lastly, according to condition (2), minimizing $\mathcal{L}_{ce}$ places the latent values for the extreme styles at the opposite ends of the canonical space. In practice, we find that training the style encoder to minimize both $\mathcal{L}_{trajectory}$ and $\mathcal{L}_{ce}$ causes the latent values of trajectories with intermediate styles to be placed monotonically between the extremes.
Summary. At training time, our architecture leverages a dataset of user demonstrations $\mathcal{D}$ and a small set of labels $\mathcal{Y}$ to learn a latent task space $Z_{\tau}$ and a canonical space of styles $Z_{\theta}$. We structure the canonical space to be consistent and monotonic by optimizing the combined loss in Equation (6). At runtime, the user selects a latent style $z_{\theta}$ from the learned canonical space. The decoder $\phi$ then takes this latent style as input and reconstructs the corresponding robot trajectory for each $z_{\tau}\in Z_{\tau}$.
@@ 5. Simulation Experiments
Environments. In the first environment we personalize the driving style of an autonomous car across two tasks, Highway and Intersection. In Highway the autonomous car follows another car on a highway. In Intersection the autonomous car waits for another car to pass before safely crossing an intersection. In both tasks, we consider 2D styles $\theta=[\theta_{1},\theta_{2}]$ that define how aggressively or defensively the car drives. Here $\theta_{1}$ corresponds to the maximum speed the car achieves in an empty section of the road, and $\theta_{2}$ represents the minimum distance that the car keeps from other vehicles on the road. For example, some users may prefer a high speed of $\theta_{1}=100$ km/h until their car is within $\theta_{2}=30$ feet of the next car. Other users may prefer a slow speed of $\theta_{1}=40$ km/h but get as close as $\theta_{2}=10$ feet of the next car. We implement these tasks using the CARLO simulator (Cao et al., 2020).
In the robot environment we move away from the conventional meaning of tasks and styles to showcase the versatility of our approach. We consider three different robot platforms as the tasks: a Kuka, Panda, and UR5. In each task (i.e., for each type of robot) the goal is to transfer a cereal box from one bin to another. The styles are three dimensional $\theta=[\theta_{1},\theta_{2},\theta_{3}]$, and represent variations in environment. The variable $\theta_{1}$ is the orientation of the cereal box, while $\theta_{2}$ and $\theta_{3}$ mark the position of the target bin along the $x$ and $z$ axis respectively. We implement this environment using Robosuite (Zhu et al., 2020).
@@
For Ours-X, all trajectories in a task are randomly sampled. We train all approaches, except Ours-L, with the same amount of labeled data. SeGMA requires labels for trajectories that belong to the same task. In contrast, Ours-X and PECAN do not require any task labels and only use labels for trajectories with similar styles. Our code can be found here: https://github.com/VT-Collab/PECAN.
Figure 3. Simulation results for autonomous driving (Top row) and robot manipulation (Bottom row) environments. We compare our proposed approach, PECAN, to a state-of-the-art baseline, SeGMA, and ablations of our approach, Ours-L and Ours-X. While SeGMA uses task labels, PECAN uses labels for trajectories with similar styles (specifically the style extremes). Both ablations use the same architecture as PECAN, however, Ours-L does not train with any labels, whereas Ours-X uses labels for trajectories with intermediate styles (instead of the style extremes). In both environments, PECAN achieves comparable Task Accuracy to SeGMA. Although PECAN has significantly lower Trajectory Error in the robot environment, its performance is similar to the baselines in the driving environment. The main advantage of PECAN over the baseline methods is that the canonical spaces learned by our approach are more consistent and monotonic (i.e., more user friendly). An asterisk (*) denotes statistical significance.
@@
Here we do not compute Consistency directly using Equation (2) because it requires access to a function $\theta(\cdot)$ that maps reconstructed trajectories to their actual style values. In our simulations, we only know the actual tasks and styles for the trajectories in the dataset $\Xi$, and not for their reconstructions. Lastly, we measure the Monotonicity of the canonical spaces by computing the correlation between the latent values and the actual styles of trajectories in $\Xi$. In a monotonous space, the difference in latent values of trajectories should be correlated to the difference in their styles. We measure this using Spearman’s rank correlation coefficient (Zar, 2005). A coefficient of $1$ or $-1$ indicates perfect correlation, while 00 means that the styles and their latent values are uncorrelated. We take the absolute value of this coefficient as the Monotonicity of the canonical space.
Results. Our results are displayed in Figure 3. We calculated these results over $20$ training runs, each starting with randomly initialized network weights. We performed one-way ANOVA tests and found that the choice of method had a significant effect on Task Accuracy in the autonomous driving ($F(3,76)=58.7$, $p<0.01$) and robot manipulation ($F(3,76)=313.1$, $p<0.01$) environments. In both environments, SeGMA achieved a high Task Accuracy, while Ours-L achieved the lowest. This is likely because SeGMA is trained with labels for the tasks, whereas Ours-L operates without any labels. PECAN, on the other hand, does not use task labels like SeGMA. Yet it achieved a comparable Task Accuracy by leveraging labels for trajectories with similar styles. Post-hoc comparisons indicated a statistically significant difference ($p<0.01$) between the Task Accuracy of PECAN and Ours-L in both environments. On the other hand, there was no significant difference in the Task Accuracy of PECAN and SeGMA in the autonomous driving ($p=0.57$) and robot manipulation ($p=0.64$) environments.
@@
Next, we observed that the canonical spaces learned by PECAN are more consistent and monotonous than those learned by any of the baselines. One-way ANOVA tests revealed that the choice of method had a significant effect on the Consistency ($F(3,76)=27.5$, $p<0.01$) and Monotonicity ($F(3,76)=30.3$, $p<0.01$) of the canonical space in the driving environment. Post-hoc comparisons indicated that the spaces learned by Ours-L show comparable Monotonicity ($p=0.53$) to PECAN but lack Consistency ($p<0.01$) because of not using any labels. In contrast, Ours-X manages to learn a consistent latent space by using the style labels similar to PECAN ($p=0.99$) in the driving environment. However, it lacks Monotonicity ($p<0.01$) because it obtains labels for the intermediate styles rather than the style extremes. SeGMA does not leverage any style labels and thus has a lower consistency ($p<0.01$) and monotonicity ($p<0.01$) than PECAN.

### v2

@@ Abstract.
Robots should personalize how they perform tasks to match the needs of individual human users. Today’s robots achieve this personalization by asking for the human’s feedback in the task space. For example, an autonomous car might show the human two different ways to decelerate at stoplights, and ask the human which of these motions they prefer. This current approach to personalization is indirect: based on the behaviors the human selects (e.g., decelerating slowly), the robot tries to infer their underlying preference (e.g., defensive driving). By contrast, our paper develops a learning and interface-based approach that enables humans to directly indicate their desired style. We do this by learning an abstract, low-dimensional, and continuous canonical space from human demonstration data. Each point in the canonical space corresponds to a different style (e.g., defensive or aggressive driving), and users can directly personalize the robot’s behavior by simply clicking on a point. Given the human’s selection, the robot then decodes this canonical style across each task in the dataset — e.g., if the human selects a defensive style, the autonomous car personalizes its behavior to drive defensively when decelerating, passing other cars, or merging onto highways. We refer to our resulting approach as PECAN: Personalizing Robot Behaviors through a Learned Canonical Space. Our simulations and user studies suggest that humans prefer using PECAN to directly personalize robot behavior (particularly when those users become familiar with PECAN), and that users find the learned canonical space to be intuitive and consistent. See videos here: https://youtu.be/wRJpyr23PKI
Representation learning, personalization, algorithmic human-robot interaction
@@
copyright: noneccs: Computing methodologies Learning latent representationsccs: Human-centered computing Graphical user interfacesccs: Computing methodologies Semi-supervised learning settings
@@ 2. Related Work
Preference Learning. Humans can personalize a robot by providing feedback about its behaviors. This includes demonstrating the desired motion (Nikolaidis et al., 2015; Ravichandar et al., 2020), correcting the robot’s actions (Li et al., 2021; Spencer et al., 2022; Losey et al., 2022), selecting their preferred trajectory from options presented by the robot (Sadigh et al., 2017; Wilde et al., 2020; Zhan et al., 2021), or some combination of these feedback types (Jain et al., 2015; Munzer et al., 2017; Bıyık et al., 2022a; Mehta and Losey, 2023). These works parameterize the robot’s behavior with user-defined features, such as the speed of an autonomous car or its distance from other cars in a driving scenario. The parameters or weights for these features are then estimated based on the human’s feedback. Together, the weights and features produce the user’s preferred behavior. We refer to this resulting behavior as the robot’s style (i.e., how it performs the given task) (Rosbach et al., 2019).
Specifying the correct set of features for modeling complex styles can be challenging (Bobu et al., 2024). While recent approaches have explored learning the features online (Katz et al., 2021), humans still need to teach new features by providing further demonstrations (Bobu et al., 2022). To overcome this, previous research has also proposed deep reinforcement learning from human feedback (Christiano et al., 2017; Hejna III and Sadigh, 2023), which bypasses the need to specify such features. However, this approach is often time-consuming, as users may have to provide feedback hundreds or even thousands of times throughout the learning process.
@@
Learning a Canonical Style Space. Most relevant to our approach are methods like (Lynch et al., 2020; Osa and Ikemoto, 2020; Jiao et al., 2022; Xihan et al., 2022) that learn a canonical representation of robot skills or styles over several tasks.
@@
In (Lynch et al., 2020), the robot embeds action sequences into a latent space of task-agnostic skills. The robot then executes these skills on a continuous range of tasks that are specified by their start and end states. When the set of tasks is finite, (Osa and Ikemoto, 2020) learns a latent space comprised of discrete and continuous portions. The discrete variables capture the tasks and the continuous variables capture the latent styles. Similar to other works in embedding robot trajectories (Co-Reyes et al., 2018; Allshire et al., 2021; Sun et al., 2021; Wang et al., 2022), both these methods employ Variational Autoencoders (VAEs) to train the latent spaces in an unsupervised manner. However, (Lynch et al., 2020) assumes that the tasks are specified by the user, while (Osa and Ikemoto, 2020) does not guarantee that the respective latent spaces will exclusively capture the tasks and styles present in the data.
@@
To address this, (Jiao et al., 2022) uses labels that specify the intention (i.e., task) and aggressiveness (i.e., style) of an autonomous car’s trajectories. Recent work (Schrum et al., 2024) has also explored adjusting the aggressiveness of autonomous driving by obtaining subjective ratings of the driving style to learn a tunable latent dimension. However, it is often challenging for humans to precisely quantify the robot’s style (Koppol et al., 2021). For this reason, (Xihan et al., 2022) instead requires users to label trajectories that belong to the same task and trajectories that correspond to the same skill. These labels are used to train Gated VAEs (Vowels et al., 2020) which encode the task and skill knowledge into separate latent spaces. A major limitation of all these methods is that the encoded styles are not consistent across tasks. The same latent value can produce different styles depending on the task — making the interface unintuitive for humans. Ideally, users should be able to select their preferred style with a single click and produce similar robot behaviors across all tasks.
Our problem also bears similarities with the style transfer problem in computer vision (Dupont, 2018; Karras et al., 2019; Jing et al., 2019; Abdal et al., 2019; Śmieja et al., 2020; Rangwani et al., 2023). For instance, (Karras et al., 2019) and (Abdal et al., 2019) transfer styles such as hair color and facial expressions from one image to another. However, this means that they require a reference style to copy from, which in our case would be a reference robot trajectory that users would have to demonstrate. In contrast, (Śmieja et al., 2020) and (Rangwani et al., 2023) generate styled images by taking a latent class and a latent style value as input. Both approaches encode classes (i.e., tasks) and styles within a single continuous latent space. Specifically, (Śmieja et al., 2020) learns a Gaussian mixture representation where each Gaussian represents a class and its variance captures the styles. Most similar to ours is Joint VAE (Dupont, 2018), which learns separate latent spaces for tasks and styles, but in an unsupervised manner. In this work, we evaluate whether these approaches apply to robotics tasks, comparing our method to (Śmieja et al., 2020) and (Dupont, 2018).
@@
Overall, similar to approaches like (Osa and Ikemoto, 2020), we utilize separate discrete and continuous latent spaces to encode the task and style information. The latent style space becomes our canonical space. To ensure that the canonical space is consistent across tasks, we use a small set of labels for trajectories having similar styles but in different tasks. Unlike (Xihan et al., 2022) and (Śmieja et al., 2020), our approach does not require any task labels. Instead, we capture the actual tasks and styles with their respective latent spaces by only using labels for trajectories with similar styles.
@@ 4.1. Separately Encoding Tasks and Styles
When this loss is minimized, it indirectly encourages the task and style encoders to capture sufficient representations of the trajectories in the dataset.
Implementation. In our experiments, we model all networks as multilayer perceptions (MLPs) with four fully connected layers and rectified linear unit (ReLU) and hyperbolic tangent (Tanh) activation functions. We flatten each trajectory into a vector with the same size as the first layer of the encoder MLPs and the last layer of the decoder MLP. For tasks with high-dimensional states, we first downsample the trajectories to reduce the size of the flattened trajectory. The task and style encoders receive the flattened trajectory as input and map it to their respective latent spaces. The latent task and style vectors are then appended and input to the trajectory decoder which outputs the flattened trajectory. We train the networks using the Adam optimizer at a learning rate of 0.0008. Our code can be found here: https://github.com/VT-Collab/PECAN
Minimizing $\mathcal{L}_{trajectory}$ trains the networks to accurately reconstruct the trajectories in our dataset. However, this still does not guarantee that the representation captured in $Z{\tau}$ aligns with the actual tasks $\mathcal{T}$, or — along the same lines — that the representation in $Z_{\theta}$ aligns with the styles $\theta$. To address this, we propose using a small set of labels $\mathcal{Y}$ for trajectories in the dataset. These labels identify trajectories from different tasks that share similar styles. In Section 4.3, we will introduce an additional loss function that leverages these labels to ensure that the representation captured in $Z_{\theta}$ aligns with the styles $\theta$. Further, we will show that by minimizing this loss together with $\mathcal{L}_{trajectory}$, we can also ensure that the latent tasks $z_{\tau}$ align with the actual tasks $\tau$.
@@ 4.2. Characteristics of a User-Friendly Canonical Space
    (2) $\theta(\phi(z_{\tau},z_{\theta}))\approx\theta(\phi(z_{\tau}^{\prime},z_{\theta}))\qquad\forall\;z_{\tau},z_{\tau}^{\prime}\in Z_{\tau}$
@@
  • Monotonicity. The styles should vary monotonically as we move from one point in the latent space to another. For example, decreasing the latent style value from $z_{\theta}=+1$ to $z_{\theta}=-1$ should gradually change the robot’s style from straight-line paths to increasingly exaggerated arm motions. Therefore, for one dimensional styles:
    (3) $(z_{\theta}-z_{\theta}^{\prime})(\theta(\phi(z_{\tau},z_{\theta}))-\theta(\phi(z_{\tau},z_{\theta}^{\prime})))\geq 0\qquad\forall\;z_{\theta},z_{\theta}^{\prime}\in Z_{\theta}$
@@ 4.3. Semi-supervised Learning
Labeling Style Extremes. We first obtain a small set of labels $\mathcal{Y}$ for trajectories having similar styles but from different tasks. Specifically, we only obtain labels for trajectories that represent the extremes of the style range. For the example in Figure 2, users may label the most exaggerated trajectories as $y_{1}$ and the least exaggerated arm motions as $y_{2}$. We believe that labeling the extremes is easier than labeling intermediate styles. For instance, users find it difficult to distinguish between slightly different arm motions (Bıyık et al., 2019). Note that we do not ask users to specify the actual style of a trajectory. We only assume that trajectories with the same label have similar styles, and correspond to one of the extremes of the canonical space.
@@
We consider each $y_{i}\in\mathcal{Y}$ to be a separate class with one-hot labels $c(y_{i})$, where $\mathcal{Y}$ contains $m$ classes. For instance, in the above example with labels $\mathcal{Y}=\{y_{1},y_{2}\}$, we would have $m=2$ classes, e.g., $c(y_{1})=[1,0]$ and $c(y_{2})=[0,1]$. At training time, we pass the labeled trajectories $\xi_{j}\in y_{i}$ through the style encoder $\psi_{\theta}$ to obtain their latent styles $z_{\theta,j}$. We then feed the latent styles into a classifier network $\Delta$ that maps each latent value to a $m$-dimensional vector of class probabilities $p_{j}=[p_{1},\ldots,p_{m}]$ such that $\sum_{k=1}^{m}p_{j,k}=1$ for any $p_{j}\in\mathcal{P}$.
@@
Lastly, according to condition (2), minimizing $\mathcal{L}_{ce}$ places the latent values for the extreme styles at the opposite ends of the canonical space. In practice, we find that training the style encoder to minimize both $\mathcal{L}_{trajectory}$ and $\mathcal{L}_{ce}$ causes the latent values of trajectories with intermediate styles to be placed monotonically between the extremes. We use equal weights for both losses when optimizing the combined loss in our experiments.
Summary. At training time, our architecture leverages a dataset of user demonstrations $\mathcal{D}$ and a small set of labels $\mathcal{Y}$ to learn a latent task space $Z_{\tau}$ and a canonical space of styles $Z_{\theta}$. We structure the canonical space to be consistent and monotonic by optimizing the combined loss in Equation (6). At runtime, the user can select any point $z_{\theta}$ inside the learned canonical space. The decoder $\phi$ then takes this latent style as input and reconstructs the corresponding robot trajectory for each $z_{\tau}\in Z_{\tau}$.
@@ 5. Simulation Experiments
Environments. In the first environment we personalize the driving style of an autonomous car across two tasks, Highway and Intersection. In Highway the autonomous car follows another car on a highway. In Intersection the autonomous car waits for another car to pass before safely crossing an intersection. The states include the 2D positions of both cars, $s=[x_{auto},y_{auto},x_{other},y_{other}]$, and the actions are the autonomous car’s velocity, $a=[\Delta x_{auto},\Delta y_{auto}].$ In both tasks, we consider 2D styles $\theta=[\theta_{1},\theta_{2}]$ that define how aggressively or defensively the car drives. Here $\theta_{1}$ corresponds to the maximum speed the car achieves in an empty section of the road, and $\theta_{2}$ represents the minimum distance that the car keeps from other vehicles on the road. For example, some users may prefer a high speed of $\theta_{1}=100$ km/h until their car is within $\theta_{2}=30$ feet of the next car. Other users may prefer a slow speed of $\theta_{1}=40$ km/h but get as close as $\theta_{2}=10$ feet of the next car. We implement these tasks using the CARLO simulator (Cao et al., 2020).
In the robot environment we move away from the conventional meaning of tasks and styles to showcase the versatility of our approach. We consider three different robot platforms as the tasks: a Kuka, Panda, and UR5. In each task (i.e., for each type of robot) the goal is to transfer a cereal box from one bin to another. The states are the joint positions and gripper configurations of the respective robot platforms and the actions are their joint and gripper velocities. Since the arms and grippers of each robot have different degrees of freedom (DoFs) — the Kuka has a $7$ DoF arm and a $6$ DoF gripper, the Panda also has a $7$ DoF arm but a $2$ DoF gripper, while the UR5 has a $6$ DoF arm and $6$ DoF gripper — we append zeros to the states and actions of the Panda and UR5 robots to ensure that all trajectories have the same number of dimensions. The styles are three dimensional $\theta=[\theta_{1},\theta_{2},\theta_{3}]$, and represent variations in environment. The variable $\theta_{1}$ is the orientation of the cereal box, while $\theta_{2}$ and $\theta_{3}$ mark the position of the target bin along the $x$ and $z$ axis respectively. We implement this environment using Robosuite (Zhu et al., 2020).
@@
We assume the dimensionality of the styles is known and model the latent style space for all methods to have the same number of dimensions as the true styles. We use a $d_{\theta}=2$ dimensional latent space for the driving environment and a $d_{\theta}=3$ dimensional space for the robot environment. Likewise, we set the number of latent tasks to match the number of tasks in each environment.
@@
For Ours-X, all trajectories in a task are randomly sampled. We train all methods using the same number of demonstrations $\mathcal{D}$ and labels $\mathcal{Y}$, except for Ours-L, which does not use the labels and only learns to reconstruct the demonstrations. SeGMA requires labels for trajectories that belong to the same task. In contrast, Ours-X and PECAN do not require any task labels and only use labels for trajectories with similar styles. These labels are provided by the expert user.
Figure 3. Simulation results for autonomous driving (Top row) and robot manipulation (Bottom row) environments. We compare our proposed approach, PECAN, to a state-of-the-art baseline, SeGMA, and ablations of our approach, Ours-L and Ours-X. While SeGMA uses task labels, PECAN uses labels for trajectories with similar styles (specifically the style extremes). Both ablations use the same architecture as PECAN, however, Ours-L does not train with any labels, whereas Ours-X uses labels for trajectories with intermediate styles (instead of the style extremes). In both environments, PECAN achieves comparable Task Accuracy to SeGMA. Although PECAN has significantly lower Trajectory Error in the robot environment, its performance is similar to the baselines in the driving environment. The main advantage of PECAN over the baseline methods is that the canonical spaces learned by our approach are more consistent and monotonic (i.e., more user friendly). An asterisk (*) denotes statistical significance and the error bars indicate standard error.
@@
Here we do not compute Consistency directly using Equation (2) because it requires access to a function $\theta(\cdot)$ that maps reconstructed trajectories to their actual style values. In our simulations, we only know the actual tasks and styles for the trajectories in the dataset $\Xi$, but not for their reconstructions. Lastly, we measure the Monotonicity of the canonical spaces by computing the correlation between the latent values and the actual styles of trajectories in $\Xi$. In a monotonous space, the difference in latent values of trajectories $\Delta Z_{\theta}$ should be rank correlated to the difference in their styles $\Delta\theta$. We measure this by calculating the Spearman’s rank correlation coefficient (Zar, 2005) between the ranks of $\Delta Z_{\theta}$ and $\Delta\theta$.
(8) $\Delta\theta=[||\theta(\xi_{1})-\theta(\xi_{2})||\;\;|\;\;\forall\,\xi_{1},\xi_{2}\in\Xi]$
$\Delta Z_{\theta}=[||\psi_{\theta}(\xi_{1})-\psi_{\theta}(\xi_{2})||\;\;|\;\;\forall\,\xi_{1},\xi_{2}\in\Xi]$
A coefficient of $1$ or $-1$ indicates perfect correlation, while 00 means that the styles and their latent values are uncorrelated. We take the absolute value of this coefficient as the Monotonicity of the canonical space.
Results. Our results are displayed in Figure 3. We calculated these results over $20$ training runs, each starting with randomly initialized network weights.
We performed one-way ANOVA tests and found that the choice of method had a significant effect on Task Accuracy in the autonomous driving ($F(3,76)=58.7$, $p<0.01$) and robot manipulation ($F(3,76)=313.1$, $p<0.01$) environments. In both environments, SeGMA achieved a high Task Accuracy, while Ours-L achieved the lowest. This is likely because SeGMA is trained with labels for the tasks, whereas Ours-L operates without any labels. PECAN, on the other hand, does not use task labels like SeGMA. Yet it achieved a comparable Task Accuracy by leveraging labels for trajectories with similar styles. Pairwise comparisons using Tukey’s HSD post-hoc test indicated a statistically significant difference ($p<0.01$) between the Task Accuracy of PECAN and Ours-L in both environments. On the other hand, there was no significant difference in the Task Accuracy of PECAN and SeGMA in the autonomous driving ($p=0.57$) and robot manipulation ($p=0.64$) environments. The p-values have been adjusted for multiple pairwise comparisons.
@@
Although Ours-L struggled to encode the tasks, it had the lowest error in reconstructing the trajectories in the driving environment. Unlike other methods that needed to balance trajectory reconstruction with shaping the canonical space, Ours-L solely focused on minimizing the trajectory loss. A one-way ANOVA test revealed a significant effect of the choice of method on Trajectory Error ($F(3,76)=4.7$, $p<0.05$) in the driving environment. However, despite of achieving the lowest trajectory error, a Tukey’s HSD post-hoc test did not indicate a significant difference ($p=0.2$) between Ours-L and PECAN. On the contrary, we found that Ours-L had significantly higher Trajectory Error ($p<0.01$) than PECAN in the robot environment. We realized that Ours-L tended to overfit to the training demonstrations in this environment leading to a poor test performance.
Finally, we observed that the canonical spaces learned by PECAN are more consistent and monotonous than those learned by any of the baselines. One-way ANOVA tests revealed that the choice of method had a significant effect on the Consistency ($F(3,76)=27.5$, $p<0.01$) and Monotonicity ($F(3,76)=30.3$, $p<0.01$) of the canonical space in the driving environment. A Tukey’s HSD post-hoc test indicated that the spaces learned by Ours-L show comparable Monotonicity ($p=0.53$) to PECAN but lack Consistency ($p<0.01$) because of not using any labels. In contrast, Ours-X manages to learn a consistent latent space by using the style labels similar to PECAN ($p=0.99$) in the driving environment. However, it lacks Monotonicity ($p<0.01$) because it obtains labels for the intermediate styles rather than the style extremes. SeGMA does not leverage any style labels and thus has a lower consistency ($p<0.01$) and monotonicity ($p<0.01$) than PECAN.